An airplane flying horizontally at a velocity of 200 m/s drops a bomb 

from an elevation of 2,415 meters. Determine the time required for the 

bomb to reach the ground


1
Expert's answer
2021-04-07T10:09:56-0400
"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot2415\ m}{9.8\ \dfrac{m}{s^2}}}=22.2\ s."